how do you solve this
x+y=9
-2x+y=-3

Answer by rfer(16322)

x+y=9, y=-x+9
-2x+y=-3
-2x-x+9=-3
-3x=-12
x=4
-----------------
4+y=9
y=5
---------------
-2*4+5=-3
-8+5=-3